# SHIELD: Secure Hypernetworks for Incremental Expansion Learning Defense

### v2

@@
Abstract
@@ 1 Introduction
Deep neural networks have achieved remarkable performance across a wide range of tasks, often surpassing human-level accuracy (He et al. 2016). However, their success hinges on static training settings that do not reflect how intelligent systems are expected to operate in the real world. In practice, models must continually adapt to changing conditions while retaining prior knowledge, a challenge known as continual learning (McCloskey and Cohen 1989; Hsu et al. 2018). At the same time, these models remain highly vulnerable to adversarial perturbations; small, carefully crafted input changes that can cause confident but incorrect predictions (Szegedy et al. 2013; Goodfellow, Shlens, and Szegedy 2014). This vulnerability raises serious concerns in high-stakes applications such as autonomous driving, robotics, and healthcare, where both adaptability and robustness are essential for safety and trust.
Surprisingly, most research treats these challenges in isolation. This disconnect is striking, given that real-world AI systems must not only learn and adapt continuously but also resist adversarial manipulation over time. Consider an autonomous vehicle that incrementally learns to navigate unfamiliar roads or environments. As it learns on the go, it must remain robust against adversarial threats like deceptive road signs, tampered traffic signals, and sensor manipulations that aim to mislead it. Without such resilience, this adaptive learning could become a critical vulnerability instead of an asset.
Despite its importance, this intersection of continual learning and adversarial robustness remains highly underexplored. Existing approaches that attempt to bridge this gap are limited. AIR (Zhou and Hua 2024), for example, uses unsupervised data augmentations to promote robustness during continual learning, but it is only demonstrated on short task sequences and cannot scale to more complex benchmarks. Double Gradient Projection (DGP) (Ru et al. 2024) enhances robustness by constraining weight updates; however, it necessitates storing large amounts of per-task gradient information. This makes it impractical for large-scale datasets or modern architectures. Moreover, this storage requirement introduces privacy concerns, especially in domains where task data is sensitive and cannot be retained.
To address these limitations, we introduce a new framework for adversarially robust continual learning. Our method uses a shared hypernetwork to generate task-specific model parameters from compact, trainable embeddings. This architecture enables efficient adaptation to new tasks without requiring replay buffers, gradient logs, or full model snapshots, thus ensuring scalability and privacy. To defend against adversarial inputs, we integrate Interval Bound Propagation (IBP) (Gowal et al. 2018), which provides formal robustness guarantees by accounting for worst-case input perturbations. We further introduce Interval MixUp, a novel technique that extends traditional MixUp by interpolating not individual input samples, but uncertainty-aware regions in the input space, enhancing robustness during training. The standard MixUp strategy is unsuitable in our setting, as training on samples with excessively large perturbation radii degrades model performance. To mitigate this, we constrain the hypercube volume near decision boundaries and employ a weighted cross-entropy loss. Empirical results demonstrate that Interval MixUp is essential for achieving high-performing models.
@@
Figure 1: SHIELD uses a hypernetwork to map task-specific embeddings $\boldsymbol{e}_{t}$ into target models that propagate virtual Interval MixUp hypercubes, enabling robust multi-task learning and certified adversarial resistance. In the Interval MixUp visualization, edge images are perturbed, the center is unperturbed, and the IBP column shows how the input hypercube is transformed across network layers.
Together, these components form a unified and efficient framework for continual learning in the presence of adversarial threats. By jointly addressing the challenges of continual adaptation and certifiable robustness, our approach takes a significant step toward building safe, trustworthy, and lifelong learning AI systems capable of operating reliably in complex, dynamic environments. An overview of our method is illustrated in Figure 1.
@@
  • SHIELD is attack-agnostic and achieves state-of-the-art or highly competitive results on multiple continual learning benchmarks, including Split miniImageNet, with up to 2$\times$ higher adversarial accuracy than prior methods.
@@ 2 Related Works
Continual learning and adversarial robustness have largely been studied in isolation, with only a few attempts to bridge the two (see Appendix A for an in-depth overview). Recent work has begun to expose the vulnerabilities of continual learning models to adversarial attacks. For example, (Khan et al. 2022) evaluates several continual learning methods under adversarial scenarios, highlighting the need for joint treatment of robustness and forgetting. AIR (Zhou and Hua 2024) offers one such integrated approach using unsupervised data augmentations for adversarial defense, though its effectiveness is limited to short task sequences.
Gradient-based methods like GEM (Lopez-Paz and Ranzato 2017) and GPM (Saha, Garg, and Roy 2021) primarily aim to prevent forgetting by using episodic memory and subspace projections, respectively, but do not explicitly address adversarial robustness. DGP (Ru et al. 2024), built as an extension of GPM, further refines these projections to maintain stability across tasks. Interestingly, the authors of DGP observe that while their method improves continual learning performance, the projection strategy itself makes the model more susceptible to adversarial attacks, revealing a critical trade-off between preserving past knowledge and ensuring robustness.
Crucially, none of these methods explore hypernetworks (Ha, Dai, and Le 2016), despite their proven benefits in continual learning (Henning et al. 2021; Książek and Spurek 2023; Wang et al. 2024) for achieving modular task representations.
@@ 3 SHIELD: Secure Hypernetworks for Incremental Expansion Learning Defense
SHIELD Architecture
Hypernetwork
Our model uses trainable embeddings $\boldsymbol{e}_{t}\in\mathbb{R}^{N},\mbox{ where }t\in\{1,...,T\}$ represents consecutive tasks indices and $T$ is its total number. These vectors serve as the input to the hypernetwork, which produces the task-dedicated target network weights. More specifically, the hypernetwork $\mathcal{H}$ with weights $\boldsymbol{\Phi}$ generates weights $\boldsymbol{\theta}_{t}$ of the target network $f$, designed for the $t$–th task, i.e. $\mathcal{H}(\boldsymbol{e}_{t};\boldsymbol{\Phi})=\boldsymbol{\theta}_{t}$. Therefore, the target network is not trained directly, and the meta-architecture generates distinct weights for each continual learning task. The function $f_{\boldsymbol{\theta}_{t}}:X\to Y$ predicts labels $Y$ for data samples $X$ based on weights $\boldsymbol{\theta}_{t}$. Finally, each continual learning task is represented by a classifier function
@@
Target network
In SHIELD, the target network is designed for interval representation of input points instead of single real values. In practice, we use the IBP approach (Gowal et al. 2018; Morawiecki et al. 2019) in the target model. Instead of utilizing a single-valued point $x=\left(x_{1},\ldots,x_{d}\right)\in\mathbb{R}^{d}$ as input to the target model $f_{\boldsymbol{\theta}_{t}}$, we propagate a $d$-dimensional hypercube:
@@
In practice, the transformations $h_{k}$ are connected by non-linear, non-decreasing activation functions such as ReLU or sigmoid, so that each layer consists of an affine mapping followed by a non-linearity. For clarity of exposition, we express each $h_{k}$ in affine form only in Eq. (3). . In the IBP approach (Gowal et al. 2018), in each layer, we find the smallest bounding box that encloses the transformed output of the previous layer. To bound the activation $z_{k}$ of the $k$–th layer in a neural network, we compute an axis-aligned hypercube $[\underline{z}_{k},\bar{z}_{k}]$, where $\underline{z}_{k}$ and $\bar{z}_{k}$ represent the element-wise lower and upper bounds, respectively. This hypercube can be written as:
@@
where $d_{k}$ is the output dimension of the $k$–th layer.
To propagate such intervals efficiently through the layers of a neural network, we use interval arithmetic, leveraging a midpoint-radius representation. This strategy enables efficient and scalable computations. Specifically, by applying Equation (3) to the input hypercube $[\underline{z}_{k-1},\bar{z}_{k-1}]$, we define:
@@
where ||$\cdot$|| ⋅ | denotes the element-wise absolute value operator. The resulting hypercube $[\underline{z}^{\star}_{k},\bar{z}^{\star}_{k}]$ serves as the input to the next activation function. For a monotonic non-decreasing activation function $g(\cdot)$, the final output bounds after applying the activation are given by:
@@
To ensure robust classification, we adopt a worst-case loss function during training (see Equation (8)) that accounts for the entire interval bound $[\underline{z}_{K},\overline{z}_{K}]$ of the final logits. This encourages the model to make consistent predictions for all inputs within a certified neighborhood of the original sample.
@@
SHIELD Loss Function
One of the core components of our model is the loss function, which simultaneously mitigates catastrophic forgetting and enhances robustness against adversarial attacks. We adopt a worst-case loss computed over the interval bounds $[\underline{z}_{K},\overline{z}_{K}]$ of the final logits. This approach ensures that the entire output bounding box is classified correctly, meaning that all permissible input perturbations within the specified interval do not alter the predicted class.
@@
As shown in (Gowal et al. 2018), computing these interval bounds requires only two forward passes through the network, making the method computationally practical. To address the challenge of overly loose bounds in complex networks, the authors of (Gowal et al. 2018) proposed a combined loss that blends the standard cross-entropy loss on the nominal inputs with the interval-based worst-case loss:
@@
where $\mathcal{L}_{\text{CE}}$ denotes the cross-entropy loss, $\sigma\left(\cdot\right)$ is a softmax function, and $\kappa\in[0,1]$ balances the trade-off. For the sake of simplicity, we omit arguments of $\mathcal{L}_{\text{IBP}}$. Notably, when $\varepsilon=0$, the interval bounds collapse to point predictions and $\mathcal{L}_{\text{IBP}}$ reduces to the standard cross-entropy loss. To facilitate stable optimization during training, we apply a scheduled annealing of $\kappa$, detailed in Appendix G.
Although ensuring adversarial robustness is possible under worst-case accuracy conditions, it is also essential to avoid catastrophic forgetting in the hypernetwork. Consequently, we incorporate a regularization term responsible for maintaining knowledge from previous continual learning tasks, as in (von Oswald et al. 2019). Therefore, we can compare the fixed hypernetwork output (i.e. the target network weights) that was generated before learning the current task $t_{c}\in\{2,\ldots,T\}$, denoted by $\boldsymbol{\Phi}^{*}$, with the hypernetwork output after recent suggestions for its weight changes, $\boldsymbol{\Phi}+\Delta\boldsymbol{\Phi}$. Ultimately, in SHIELD, the formula for the regularization loss is presented as:
@@
Robustness Guarantees of SHIELD
@@
$\underline{z}^{(K)}_{y_{\text{true}}}\left(x;\boldsymbol{\theta}_{t}\right)>\max_{j\neq y_{\text{true}}}\overline{z}^{(K)}_{j}\left(x;\boldsymbol{\theta}_{t}\right),$ (12)
for all the samples from the perturbation set $\{x+\delta:\|\delta\|_{\infty}\leq\varepsilon\}$, where $\underline{z}^{(K)}$, $\overline{z}^{(K)}$ are lower and upper bounds, respectively, on the output logits for $x$ at the final layer $K$, computed using IBP.
@@
Consider a continual learning setup with a sequence of $T$ tasks, $\{\mathcal{D}_{t}\}_{t=1}^{T}$, where each task dataset $\mathcal{D}_{t}=\{(x_{i}^{(t)},y_{i}^{(t)})\}_{i=1}^{N_{t}}$ is drawn from a task-specific distribution $P_{t}(x,y)$. Let $\boldsymbol{\theta}_{t}$ denote the model parameters after learning task $t$. The model is said to be certifiably robust up to time $t$ if, for all samples $(x,y_{\text{true}})\in\bigcup_{s=1}^{t}\mathcal{D}_{s}$, the following condition holds:
$\underline{z}^{(K)}_{y_{\text{true}}}(x;\boldsymbol{\theta}_{t})>\max_{j\neq y_{\text{true}}}\overline{z}^{(K)}_{j}(x;\boldsymbol{\theta}_{t}),\quad\text{for all }\|\delta\|_{\infty}\leq\varepsilon.$ (13)
@@
This condition guarantees that the model’s predictions remain robust to bounded adversarial perturbations across all tasks encountered so far, without requiring access to the full datasets $\{\mathcal{D}_{t}\}_{t=1}^{T}$.
SHIELD is designed to meet this robustness criterion by integrating IBP-based training within a hypernetwork-based continual learning framework. Specifically, we verify that for any task $s\leq t$ and any sample $(x,y_{\text{true}})\in\mathcal{D}_{s}$, Equation (13) holds at any time $t$, ensuring robustness to perturbations within an $\varepsilon$-bounded $\ell_{\infty}$-norm.
In our approach, the hypernetwork generates task-specific parameters $\boldsymbol{\theta}_{t}=\mathcal{H}(\boldsymbol{e}_{t};\boldsymbol{\Phi})$ for the target network $f_{\boldsymbol{\theta}_{t}}$. We train the hypernetwork using IBP to minimize the certified loss, as outlined in Equation (9). This training procedure encourages each task-specific model to be certifiably robust within its assigned interval bounds. To mitigate forgetting and support robustness across tasks, the regularization term in Equation (10) promotes consistency with previously generated task parameters. As a result, our method supports the accumulation of certified robustness across tasks throughout continual learning.
Assuming effective optimization of both the certified loss and the regularization term, the resulting model $f_{\boldsymbol{\theta}_{t}}$ satisfies $\underline{z}^{(K)}_{y_{\text{true}}}(x;\boldsymbol{\theta}_{t})>\max_{j\neq y_{\text{true}}}\overline{z}^{(K)}_{j}(x;\boldsymbol{\theta}_{t})$, for all samples $(x,y_{\text{true}})$ from the combined datasets $\bigcup_{s=1}^{t}\mathcal{D}_{s}$ and all perturbations $\|\delta\|_{\infty}\leq\varepsilon$.

### v3

@@
SHIELD: Secure Hypernetworks for Incremental Expansion Learning Defense
Patryk Krukowski
Jagiellonian University
IDEAS NCBR
patryk.krukowski@doctoral.uj.edu.pl
   Łukasz Gorczyca
Jagiellonian University
lukasz.gorczyca@uj.edu.pl
   Piotr Helm
Jagiellonian University
piotr.helm@uj.edu.pl
   Kamil Książek
Jagiellonian University
kamil.ksiazek@uj.edu.pl
   Przemysław Spurek
Jagiellonian University
IDEAS Research Institute
przemyslaw.spurek@uj.edu.pl
Streszczenie
@@ 1 Introduction
Deep neural networks have achieved remarkable performance across a wide range of tasks, often surpassing human-level accuracy [he2016deep]. However, their success hinges on static training settings that do not reflect how intelligent systems are expected to operate in the real world. In practice, models must continually adapt to changing conditions while retaining prior knowledge, a challenge known as continual learning [mccloskey1989catastrophic, hsu2018re]. At the same time, these models remain highly vulnerable to adversarial perturbations; small, carefully crafted input changes that can cause confident but incorrect predictions [szegedy2013intriguing, goodfellow2014explaining]. This vulnerability raises serious concerns in high-stakes applications such as autonomous driving, robotics, and healthcare, where both adaptability and robustness are essential for safety and trust.
Surprisingly, most research treats these challenges in isolation. This disconnect is striking, given that real-world AI systems must not only learn and adapt continuously but also resist adversarial manipulation over time. Consider an autonomous vehicle that incrementally learns to navigate unfamiliar roads or environments. As it learns on the go, it must remain robust against adversarial threats, such as deceptive road signs, tampered traffic signals, and sensor manipulations that aim to mislead it. Without such resilience, this adaptive learning could become a critical vulnerability instead of an asset.
@@
Rysunek 1: SHIELD uses a hypernetwork to map task-specific embeddings $\boldsymbol{e}_{t}$ into target models that propagate virtual Interval MixUp hypercubes, enabling robust multi-task learning and certified adversarial resistance. In the Interval MixUp visualization, edge images are perturbed, the center is unperturbed, and the IBP column shows how the input hypercube is transformed across network layers.
Despite its importance, this intersection of continual learning and adversarial robustness remains highly underexplored. Existing approaches that attempt to bridge this gap are limited. AIR [zhou2024defense], for example, uses unsupervised data augmentations to promote robustness during continual learning, but it is only demonstrated on short task sequences and cannot scale to more complex benchmarks. Double Gradient Projection (DGP) [ru2024maintaining] enhances robustness by constraining weight updates; however, it necessitates storing large amounts of per-task gradient information. This makes it impractical for large-scale datasets or modern architectures. Moreover, this storage requirement introduces privacy concerns, especially in domains where task data is sensitive and cannot be retained.
To address these limitations, we introduce a new framework for adversarially robust continual learning. Our method uses a shared hypernetwork to generate task-specific model parameters from compact, trainable embeddings. This architecture enables efficient adaptation to new tasks without requiring replay buffers, gradient logs, or full model snapshots, thus ensuring scalability and privacy. To defend against adversarial inputs, we integrate Interval Bound Propagation (IBP) [gowal2018effectiveness], which provides formal robustness guarantees by accounting for worst-case input perturbations. We further introduce Interval MixUp, a novel technique that extends traditional MixUp by interpolating not individual input samples, but uncertainty-aware regions in the input space, enhancing robustness during training. The standard MixUp strategy is unsuitable in our setting, as training on samples with excessively large perturbation radii degrades model performance. To mitigate this, we constrain the hypercube volume near decision boundaries and employ a weighted cross-entropy loss. Empirical results demonstrate that Interval MixUp is essential for achieving high-performing models.
Together, these components form a unified and efficient framework for continual learning in the presence of adversarial threats. By jointly addressing the challenges of continual adaptation and certifiable robustness, our approach takes a significant step toward building safe, trustworthy, and lifelong learning AI systems capable of operating reliably in complex, dynamic environments. An overview of our method is illustrated in Fig. 1.
@@
[…]
@@ 2 Related Works
Continual learning and adversarial robustness have largely been studied in isolation, with only a few attempts to bridge the two (see Supplementary Material (SM) 7 for an in-depth overview). Recent work has begun to expose the vulnerabilities of continual learning models to adversarial attacks. For example, [khan2022susceptibility] evaluates several continual learning methods under adversarial scenarios, highlighting the need for joint treatment of robustness and forgetting. AIR [zhou2024defense] offers one such integrated approach using unsupervised data augmentations for adversarial defense, though its effectiveness is limited to short task sequences.
Gradient-based methods like GEM [Paz2017gradient] and GPM [saha2021gradient] primarily aim to prevent forgetting by using episodic memory and subspace projections, respectively, but do not explicitly address adversarial robustness. DGP [ru2024maintaining], built as an extension of GPM, further refines these projections to maintain stability across tasks. Interestingly, the authors of DGP observe that while their method improves continual learning performance, the projection strategy itself makes the model more susceptible to adversarial attacks, revealing a critical trade-off between preserving past knowledge and ensuring robustness.
Crucially, none of these methods explore hypernetworks [ha2016hypernetworks], despite their proven benefits in continual learning [henning2021posterior, ksikazek2023hypermask, wang2024comprehensive] for achieving modular task representations.
@@ 3 SHIELD: Secure Hypernetworks for Incremental Expansion Learning Defense
3.1 SHIELD Architecture
Hypernetwork.
Our model uses trainable embeddings $\boldsymbol{e}_{t}\in\mathbb{R}^{N},\mbox{ where }t\in\{1,...,T\}$ represents consecutive tasks indices and $T$ is its total number. These vectors serve as the input to the hypernetwork, which produces the task-dedicated target network weights. More specifically, the hypernetwork $\mathcal{H}$ with weights $\boldsymbol{\Phi}$ generates weights $\boldsymbol{\theta}_{t}$ of the target network $f$, designed for the tt–th task, i.e. $\mathcal{H}(\boldsymbol{e}_{t};\boldsymbol{\Phi})=\boldsymbol{\theta}_{t}$. Therefore, the target network is not trained directly, and the meta-architecture generates distinct weights for each continual learning task. The function $f_{\boldsymbol{\theta}_{t}}:X\to Y$ predicts labels $Y$ for data samples $X$ based on weights $\boldsymbol{\theta}_{t}$. Finally, each continual learning task is represented by a classifier function
@@ 3.1 SHIELD Architecture
Target Network.
In SHIELD, the target network is designed for interval representation of input points instead of single real values. In practice, we use the IBP approach [gowal2018effectiveness, morawiecki2019fast] in the target model. Instead of utilizing a single-valued point $x=\left(x_{1},\ldots,x_{d}\right)\in\mathbb{R}^{d}$ as input to the target model $f_{\boldsymbol{\theta}_{t}}$, we propagate a $d$-dimensional hypercube:
@@
In practice, the transformations $h_{k}$ are connected by non-linear, non-decreasing activation functions such as ReLU or sigmoid, so that each layer consists of an affine mapping followed by a non-linearity. For clarity of exposition, we express each $h_{k}$ in affine form only in Eq. (3). . In the IBP approach [gowal2018effectiveness], in each layer, we find the smallest bounding box that encloses the transformed output of the previous layer. To bound the activation $z_{k}$ of the kk–th layer in a neural network, we compute an axis-aligned hypercube $\underline{z}_{k},\bar{z}_{k}]$, where $\underline{z}_{k}$ and $\bar{z}_{k}$ represent the element-wise lower and upper bounds, respectively. This hypercube can be written as:
@@
where $d_{k}$ is the output dimension of the kk–th layer.
To propagate such intervals efficiently through the layers of a neural network, we use interval arithmetic, leveraging a midpoint-radius representation. This strategy enables efficient and scalable computations. Specifically, by applying Eq. (3) to the input hypercube $\underline{z}_{k-1},\bar{z}_{k-1}]$, we define:
@@
where $|\cdot|$ denotes the element-wise absolute value operator. The resulting hypercube $\underline{z}^{\star}_{k},\bar{z}^{\star}_{k}]$ serves as the input to the next activation function. For a monotonic non-decreasing activation function $g(\cdot)$, the final output bounds after applying the activation are given by:
@@
To ensure robust classification, we adopt a worst-case loss function during training (see Eq. (8)) that accounts for the entire interval bound $\underline{z}_{K},\overline{z}_{K}]$ of the final logits. This encourages the model to make consistent predictions for all inputs within a certified neighborhood of the original sample.
@@
3.2 SHIELD Loss Function
One of the core components of our model is the loss function, which simultaneously mitigates catastrophic forgetting and enhances robustness against adversarial attacks. We adopt a worst-case loss computed over the interval bounds $\underline{z}_{K},\overline{z}_{K}]$ of the final logits. This approach ensures that the entire output bounding box is classified correctly, meaning that all permissible input perturbations within the specified interval do not alter the predicted class.
@@ 3.2 SHIELD Loss Function
As shown in [gowal2018effectiveness], computing these interval bounds requires only two forward passes through the network, making the method computationally practical. To address the challenge of overly loose bounds in complex networks, the authors of [gowal2018effectiveness] proposed a combined loss that blends the standard cross-entropy loss on the nominal inputs with the interval-based worst-case loss:
@@
where $\mathcal{L}_{\text{CE}}$ denotes the cross-entropy loss, $\sigma\left(\cdot\right)$ is a softmax function, and $\kappa\in[0,1]$ balances the trade-off. For the sake of simplicity, we omit arguments of $\mathcal{L}_{\text{IBP}}$. Notably, when $\varepsilon=0$, the interval bounds collapse to point predictions and $\mathcal{L}_{\text{IBP}}$ reduces to the standard cross-entropy loss. To facilitate stable optimization during training, we apply a scheduled annealing of $\kappa$, detailed in SM 13. Additionally, we study the effect of different choices of $\kappa$ on the final performance in SM 14.
Although ensuring adversarial robustness is possible under worst-case accuracy conditions, it is also essential to avoid catastrophic forgetting in the hypernetwork. Consequently, we incorporate a regularization term responsible for maintaining knowledge from previous continual learning tasks, as in [von2019continual]. Therefore, we can compare the fixed hypernetwork output (i.e. the target network weights) that was generated before learning the current task $t_{c}\in\{2,\ldots,T\}$, denoted by $\boldsymbol{\Phi}^{*}$, with the hypernetwork output after recent suggestions for its weight changes, $\boldsymbol{\Phi}+\Delta\boldsymbol{\Phi}$. Ultimately, in SHIELD, the formula for the regularization loss is presented as:
@@
3.3 Robustness Guarantees of SHIELD
@@
$\underline{z}^{(K)}_{y_{\text{true}}}\left(x;\boldsymbol{\theta}\right)>\max_{j\neq y_{\text{true}}}\overline{z}^{(K)}_{j}\left(x;\boldsymbol{\theta}\right),$ (12)
where $\underline{z}^{(K)}$, $\overline{z}^{(K)}$ are lower and upper bounds, respectively, on the output logits for $x$ at the final layer $K$, computed using IBP.
@@
Consider a continual learning setup with a sequence of $T$ tasks, $\{\mathcal{D}_{t}\}_{t=1}^{T}$, where each task dataset $\mathcal{D}_{t}=\{(x_{i}^{(t)},y_{i}^{(t)})\}_{i=1}^{N_{t}}$. Let $\boldsymbol{\theta}_{t}$ denote the model parameters after learning task $t$. The model is said to be certifiably robust up to time $t$ if, for all samples $(x,y_{\text{true}})\in\bigcup_{s=1}^{t}\mathcal{D}_{s}$, the following condition holds:
$\underline{z}^{(K)}_{y_{\text{true}}}(x;\boldsymbol{\theta}_{t})>\max_{j\neq y_{\text{true}}}\overline{z}^{(K)}_{j}(x;\boldsymbol{\theta}_{t}).$ (13)
@@
This condition guarantees that the model’s predictions remain robust to bounded adversarial perturbations across all tasks encountered so far.
To understand how robustness can be preserved across tasks, we analyze the effect of parameter updates during continual learning. When the model begins learning task $(t+1)$, the parameters of the hypernetwork are updated, which in turn alters the classifier weights it generates. We denote by $\boldsymbol{\theta}_{s,t}$ the classifier weights produced by the hypernetwork to solve the $s$-th task after learning task $t$. Consequently, these weights are modified from $\boldsymbol{\theta}_{s,t}$ to $\boldsymbol{\theta}_{s,t}+\boldsymbol{h}$, where $\boldsymbol{h}$ denotes the induced update. To maintain the robustness guarantees established for any previous task $s\leq t$, this update must not violate the robustness condition given in Eq. (13).
For notational simplicity, when we write $\boldsymbol{\theta}_{t}$ with a single subscript, we refer to the classifier weights generated to solve the $t$-th task, without specifying the stage of continual training at which they were produced.
We characterize how much the classifier outputs may change under the update $\boldsymbol{h}$ while still preserving the previously certified robustness. This leads to the following sufficient condition for robustness preservation.
Theorem 3.1 (Certified Robustness Preservation).
Suppose that learning task $(t+1)$ updates the classifier parameters by $\boldsymbol{h}$, yielding $\boldsymbol{\theta}_{s,t}+\boldsymbol{h}$. For any previous task $s\in\{1,\ldots,t\}$, let the certified margin at time $t$ for a sample $(x,y_{\text{true}})\in\mathcal{D}_{s}$ be defined as $M(x,y_{\text{true}};\boldsymbol{\theta}_{s,t})=\underline{z}^{(K)}_{y_{\text{true}}}\left(x;\boldsymbol{\theta}_{s,t}\right)-\max_{j\neq y_{\text{true}}}\overline{z}^{(K)}_{j}\left(x;\boldsymbol{\theta}_{s,t}\right).$
Let $\Delta_{\text{max}}^{(s,t)}(x;\boldsymbol{h})$ denote the maximum change in the classifier’s output logits on $(x,y_{\text{true}})$ due to the parameter update $\boldsymbol{h}$: $\Delta_{\text{max}}^{(s,t)}(x;\boldsymbol{h})=\|f(x;\boldsymbol{\theta}_{s,t}+\boldsymbol{h})-f(x;\boldsymbol{\theta}_{s,t})\|_{\infty}.$ Then, a continual learner preserves certified robustness for $(x,y_{\text{true}})\in\mathcal{D}_{s}$ if the parameter update $\boldsymbol{h}$ satisfies
$\Delta{\text{max}}^{(s,t)}(x;\boldsymbol{h})\leq\tfrac{1}{2}M(x,y_{\text{true}};\boldsymbol{\theta}_{s,t}).$ (14)
Proof is provided in SM 21.
SHIELD unifies IBP-based robust training with a hypernetwork-driven continual learning framework to maintain certified robustness across sequential tasks. Specifically, the hypernetwork generates task-specific parameters $\boldsymbol{\theta}_{t}=\mathcal{H}(\boldsymbol{e}_{t};\boldsymbol{\Phi})$, while IBP training (Eq. (9)) enforces that each task model remains certifiably robust within its interval bounds. To mitigate catastrophic forgetting, the regularization term in Eq. (10) enforces consistency with parameters produced for prior tasks, stabilizing learned representations and preserving previously acquired knowledge. Importantly, as stated in Theorem 3.1, when this regularization is sufficiently strong, it ensures not only the retention of past task performance but also the preservation of their certified robustness guarantees. We additionally provide a theoretical analysis of SHIELD in SM 21, formally establishing its robustness-preserving continual learning capabilities.